# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ Appendix D Bit threads from PDE
The max-flow formulation of holographic entanglement entropy, seeking a divergenceless vector field $v^{a}$ with bounded norm $|v|\leq\mathrm{C}$ that maximises flux $\int_{A}v$, bears a structural resemblance to certain problems involving partial differential equations (PDEs). This appendix aims to briefly outline these connections.
One relevant connection arises from potential flow theory. If the bit thread field $v^{a}$ were derivable from a scalar potential $\psi$ (i.e., $v^{a}=\nabla^{a}\psi$), the divergenceless condition $\nabla_{a}v^{a}=0$ would translate into the Laplace equation $\nabla^{2}\psi=0$. The problem would then involve solving Laplace equation subject to the non-linear norm constraint $|\nabla\psi|\leq\mathrm{C}$ and appropriate boundary conditions related to the flux maximisation for region $A$.
Figure 3: Flow lines for bit threads obtained from solving the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and imposing the conditions $v_{PDE}^{a}|_{{RT}}=n^{a}$ and $|v_{PDE}|\leq 1$. The blue horizontal line denotes the boundary subregion and the red semicircle is the corresponding RT surface.
Here we construct a flow directly from the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and using the method of characteristics, we find a general solution. Then imposing the conditions $v_{PDE}^{a}|_{\footnotesize{RT}}=n^{a}/4G_{N}$ and $|v_{PDE}|\leq 1/4G_{N}$. We have not shown for the scalar $\psi$, as we are interested in the vector $\nabla^{a}\psi$ itself. For convenience, we solved the PDE in the Poincaré AdS3 background (2.19) and find a vector field given by,
@@
This is a valid bit thread configuration. Note that this vector is only well defined from the boundary subregion $A$ to the corresponding RT surface $m(A)$. Beyond that we need to glue some other thread configuration (like the one in section 2 or the geodesic ones) such that $v^{a}$ remains continuous and differentiable at the RT surface. A similar method was used for ‘maximally packed flows’ in [18]. The flow lines of $v_{PDE}^{a}$ are shown in figure 3. For simple geometries this method can also be applied to construct a flow for bit threads.
@@ Appendix E Coordinate systems of AdS3
Here we will discuss the relation between two different coordinate patches of the AdS spacetime, namely the Poincaré and Rindler AdS. We will mainly follow the notations of [27]. The geometry of AdSd+1 can be expressed by embedding it to a $\mathbb{R}^{2,d}$ space as
@@
As discussed in section 2, we are mainly interested in 3 dimensional AdS spacetime. So, the above geometry reduces to AdS3 embedded in $\mathbb{R}^{2,2}$ space as
@@
Combing the above equation with the hyperboloid (E.3) gives the standard Poincaré AdS3 metric
@@
As our main objective is to use the coordinate transformations to find a vector field in terms of one coordinate system from the other on a fixed Cauchy surface, we will look at the above relations with $t=0=\tau$ which is a common Cauchy surface in both the coordinate systems. Now, the above relations simplify to
@@
Using these coordinate transformations one can check that the bifurcation surface of (E.8) intersects the boundary of the Poincaré coordinates on a sphere of radius $L$ in the boundary CFT with flat background [27]. This sphere is the entangling region for which we want to obtain the bit threads configuration.
The above discussion of coordinate transformations can be extended for a sphere of arbitrary radius $R$ in the CFT [27]. It implies that the horizon (or more specifically the bifurcation surface) of the topological black hole moves to a different position and the intersection with the boundary also changes. In the Poincaré coordinates, the change of position of the bifurcation surface corresponds to the different RT surfaces (arcs) corresponding to the boundary entangling region. One can find a family of arcs which are connected by the isometries of the AdS geometry. A boost in the $(Y_{-1},Y_{2})$ plane of the embedding space is one such isometry. Under this boost, the coordinates transform as
@@
while keeping the other coordinates remain same. Using these we again end up with the metric (E.8) and choose the Poincaré coordinates to be same as in the (E.6). To relate these two coordinates, we need to write $Y_{-1}$ and $Y_{2}$ in terms of the boost parameter $\beta$ as

### v2

@@ Appendix D Bit threads from PDE
The max-flow formulation of holographic entanglement entropy, seeking a divergenceless vector field $v^{a}$ with bounded norm $|v|\leq\mathrm{C}$ that maximises flux $\int_{A}v$, bears a structural resemblance to specific problems involving partial differential equations (PDEs). This appendix aims to outline these connections briefly.
One relevant connection arises from potential flow theory. If the bit thread field $v^{a}$ were derivable from a scalar potential $\psi$ (i.e., $v^{a}=\nabla^{a}\psi$), the divergenceless condition $\nabla_{a}v^{a}=0$ would translate into the Laplace equation $\nabla^{2}\psi=0$. The problem would then involve solving the Laplace equation subject to the non-linear norm constraint $|\nabla\psi|\leq\mathrm{C}$ and appropriate boundary conditions related to the flux maximisation for region $A$.
Figure 3: Flow lines for bit threads obtained from solving the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and imposing the conditions $v_{PDE}^{a}|_{RT}=n^{a}$ and $|v_{PDE}|\leq 1$. The blue horizontal line denotes the boundary subregion, and the red semicircle is the corresponding RT surface.
Here, we construct a flow directly from the PDE: $\nabla_{a}v_{PDE}^{a}=0$ and, using the method of characteristics, we find a general solution. Then we impose the conditions $v_{PDE}^{a}|_{RT}=n^{a}/4G_{N}$ and $|v_{PDE}|\leq 1/4G_{N}$. For convenience, we solved the PDE in the Poincaré AdS3 background (2.19) and find a vector field given by,
@@
This is a valid bit thread configuration. Note that this vector is only well defined from the boundary subregion $A$ to the corresponding RT surface $m(A)$. Beyond that, we need to glue some other thread configuration such that $v^{a}$ remains continuous and differentiable at the RT surface. A similar method was used for ‘maximally packed flows’ in [3]. The flow lines of $v_{PDE}^{a}$ are shown in figure 3. For simple geometries, this method can also be applied to construct a flow for bit threads.
@@ Appendix E Coordinate systems of AdS3
Here we will discuss the relation between two different coordinate patches of the AdS spacetime, namely the Poincaré and Rindler AdS. We will mainly follow the notation of [17]. The geometry of AdSd+1 can be expressed by embedding it to a $\mathbb{R}^{2,d}$ space as
@@
As discussed in section 2, we are mainly interested in 3-dimensional AdS spacetime. So, the above geometry reduces to AdS3 embedded in $\mathbb{R}^{2,2}$ space as
@@
Combining the above equation with the hyperboloid (E.3) gives the standard Poincaré AdS3 metric
@@
Our main objective is to use the coordinate transformations to find a vector field in terms of one coordinate system from the other on a fixed Cauchy surface. So we will use the above relations with $t=0=\tau$, which is a common Cauchy surface in both coordinate systems. Now, the above relations simplify to
@@
Using these coordinate transformations one can check that the bifurcation surface of (E.8) intersects the boundary of the Poincaré coordinates on a sphere of radius $L$ in the boundary CFT with flat background [17]. This sphere is the entangling region for which we want to obtain the bit thread configuration.
The above discussion of coordinate transformations can be extended for a sphere of arbitrary radius $R$ in the CFT [17]. It implies that the horizon (or more specifically, the bifurcation surface) of the topological black hole moves to a different position, and the intersection with the boundary also changes. In the Poincaré coordinates, the change of position of the bifurcation surface corresponds to the different RT surfaces (arcs) corresponding to the boundary entangling region. One can find a family of arcs that are connected by the isometries of the AdS geometry. A boost in the $(Y_{-1},Y_{2})$ plane of the embedding space is one such isometry. Under this boost, the coordinates transform as
@@
while keeping the other coordinates the same. Using these, we again end up with the metric (E.8) and choose the Poincaré coordinates to be the same as in the (E.6). To relate these two coordinates, we need to write $Y_{-1}$ and $Y_{2}$ in terms of the boost parameter $\beta$ as